If you wanted to construct a 95% confidence interval for the mean (with σ being unknown), what level of significance would you use when referencing the t distributions table?

0.95

0.5

0.05

90210
Answer #1

Solution:

Given that,

At 95% confidence level the is ,

= 1 - 95% = 1 - 0.95 = 0.05

significance level = 0.05
Answer #2

Explanation:

When constructing a 95% confidence interval, the level of significance (α) is calculated as:

α=1Confidence Level=10.95=0.05

This means you would use 0.05 when referencing the t-distribution table to find the critical t-value.


source: Self
answered by: Harshwardhan kunal